OpenStudy (anonymous):

2x − 2y − 5z + 2x + 2y + 3z

OpenStudy (callisto):

2x − 2y − 5z + 2x + 2y + 3z = 2x + 2x − 2y + 2y − 5z + 3z = can you do the addition/subtraction for the like terms?

OpenStudy (anonymous):

4x-y-2z

OpenStudy (anonymous):

\[2x-2y-5z+2x+2y+3z=4x-2z\]

OpenStudy (anonymous):

Thank you for the medal